OpenStudy (anonymous):

solve x^2+4x-4=0!!! Please help. Best answer gets the medal. Pleaseeee help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! choices x equals the quantity of negative 2 plus or minus square root 16 all over 2 x equals negative 4 plus or minus square root 2 x equals negative 2 plus or minus 2 square root 2 x equals the quantity of 4 plus or minus square root 18 all over 2

OpenStudy (solomonzelman):

What do you want to do, complete the square, factor, or quadratic formula?

OpenStudy (anonymous):

quadratic form

OpenStudy (anonymous):

formula

OpenStudy (solomonzelman):

\[\huge {\color{blue} {x=\frac{-b ±\sqrt{b^2-4ac}}{2a}} }\]

OpenStudy (solomonzelman):

\[\huge {\color{blue} {ax^2+bx+c=0} }\]

OpenStudy (anonymous):

the right answer is x=-2+2sqrt2 but i need to know how to solve it

OpenStudy (solomonzelman):

see my ax^2+bx.... ? find your a=? b=? c=? plug them in.

OpenStudy (anonymous):

i have no clue can you just show me how to solve it?? please

OpenStudy (solomonzelman):

I'll give you an example, ready?

OpenStudy (anonymous):

can you just show me how to solve this one please last Q of hw

OpenStudy (solomonzelman):

OK.

OpenStudy (solomonzelman):

looking at \[\huge {\color{blue} {ax^2+bx+c=0} }\] compare to \[\huge {\color{red} {1} }\huge {\color{blue} {x^2+} }\huge {\color{red} {4} }\huge {\color{blue} {x} } \huge {\color{red } {-4} }=\huge {\color{blue} {0} }\] see the coefficients in red?

OpenStudy (solomonzelman):

So, \[\huge {\color{green} {a=1} }\]\[\huge {\color{green } {b=4} }\]\[\huge {\color{green } {c=-4} }\]

OpenStudy (solomonzelman):

I'll show fast.

OpenStudy (solomonzelman):